# Uncovering the Mechanism of Chiral Three-Nucleon Force in Driving Spin-Orbit Splitting

### v1

@@ B.0.7 Two-pion exchange $c_{4}$ term
$w_{3N}^{(c_{4})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$ $=\frac{g_{A}^{2}c_{4}}{8f_{\pi}^{4}}\frac{q_{i}^{2}q_{j}^{2}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\sum_{\lambda=0}^{3}\mathcal{O}_{\lambda}^{(c_{4})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (30)
$\mathcal{O}_{\lambda}^{(c_{4})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=4\pi i\sqrt{6}\hat{\lambda}\sum_{l_{i}l_{j}S}\hat{S}\left(1010|l_{i}0\right)\left(1010|l_{j}0\right)\begin{Bmatrix}l_{i}&l_{j}&\lambda\\ 1&1&S\\ 1&1&1\end{Bmatrix}$
×$\times\left[\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{S}\otimes\sigma_{1}(k)\right]_{\lambda}\otimes\left[Y_{l_{i}}\!\left(\hat{\bm{q}}_{i}\right)\otimes Y_{l_{j}}\!\left(\hat{\bm{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (31)
@@
$\mathcal{O}_{0}^{(c_{4})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=\frac{1}{6}\left[\left(\bm{\sigma}_{i}\times\bm{\sigma}_{j}\right)\cdot\bm{\sigma}_{k}\right]\left[1-\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)^{2}\right],$ (32)
$\mathcal{O}_{1}^{(c_{4})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=\frac{1}{10}\left[4\left(\bm{\sigma}_{i}\cdot\bm{\sigma}_{j}\right)\bm{\sigma}_{k}-\left(\bm{\sigma}_{j}\cdot\bm{\sigma}_{k}\right)\bm{\sigma}_{i}-\left(\bm{\sigma}_{k}\cdot\bm{\sigma}_{i}\right)\bm{\sigma}_{j}\right]\cdot\left(\hat{\bm{q}}_{i}\times\hat{\bm{q}}_{j}\right)\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right).$ (33)
@@
Appendix C Effective single-particle energy
@@
Table 2: Occupation numbers $\Braket{\hat{N}_{i}}$ for ${}^{8}\mathrm{Be}$ and ${}^{12}\mathrm{C}$. The column “No 3NF” is the results obtained with the 2NF only, while the other columns are those including the 3NFs of rank $\lambda$.

### v2

@@ A.6 Two-pion exchange $c_{4}$ term
$w_{3N}^{(c_{4})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$ $=\frac{g_{A}^{2}c_{4}}{8f_{\pi}^{4}}\frac{q_{i}^{2}q_{j}^{2}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}$
×$\times\sum_{\lambda=0}^{3}\mathcal{O}_{\lambda}^{(c_{4})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (30)
$\mathcal{O}_{\lambda}^{(c_{4})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=4\pi i\sqrt{6}\hat{\lambda}\sum_{l_{i}l_{j}S}\hat{S}\left(1010|l_{i}0\right)\left(1010|l_{j}0\right)\begin{Bmatrix}l_{i}&l_{j}&\lambda\\ 1&1&S\\ 1&1&1\end{Bmatrix}$
×$\quad\times\left[\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{S}\otimes\sigma_{1}(k)\right]_{\lambda}\otimes\left[Y_{l_{i}}\!\left(\hat{\boldsymbol{q}}_{i}\right)\otimes Y_{l_{j}}\!\left(\hat{\boldsymbol{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (31)
@@
$\mathcal{O}_{0}^{(c_{4})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=\frac{1}{6}\left[\left(\boldsymbol{\sigma}_{i}\times\boldsymbol{\sigma}_{j}\right)\cdot\boldsymbol{\sigma}_{k}\right]\left[1-\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)^{2}\right],$ (32)
$\mathcal{O}_{1}^{(c_{4})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=\frac{1}{10}\left[4\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{j}\right)\boldsymbol{\sigma}_{k}-\left(\boldsymbol{\sigma}_{j}\cdot\boldsymbol{\sigma}_{k}\right)\boldsymbol{\sigma}_{i}-\left(\boldsymbol{\sigma}_{k}\cdot\boldsymbol{\sigma}_{i}\right)\boldsymbol{\sigma}_{j}\right]\cdot\left(\hat{\boldsymbol{q}}_{i}\times\hat{\boldsymbol{q}}_{j}\right)$
×$\quad\times\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right).$ (33)
@@
Appendix B Effective single-particle energy
@@
Table 2: Occupation numbers $\Braket{\hat{N}_{i}}$ for ${}^{8}\mathrm{Be}$ and ${}^{12}\mathrm{C}$. The columns denoted by “No 3NF” are the results obtained with the 2NF only, while the other columns are those including the 3NFs of rank $\lambda$.
@@
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.